For each of the following strong base solutions, determine [OH−],[H3O+], pH, and pOH.

A) 1.6×10−3 M Ca(OH)2

Express your answer using two significant figures. (OH-), (H3O+)

Enter your answers numerically separated by commas. pH, pOH. Express your answer to two decimal places. Enter your answers numerically separated by commas.

B) 4.9×10−4 M Sr(OH)2

Express your answer using two significant figures. Enter your answers numerically separated by commas: pH, pOH. Express your answer to two decimal places. Enter your answers numerically separated by commas.

C) 8.7×10−5 M KOH

Express your answer using two significant figures. Enter your answers numerically separated by commas: pH, pOH. Express your answer to two decimal places. Enter your answers numerically separated by commas.
Answer #1

A)

[OH-] = 2*[Ca(OH)2]

= 2*1.6*10^-3 M

= 3.2*10^-3 M

use:

[H+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H+] = (1.0*10^-14)/[OH-]

[H+] = (1.0*10^-14)/3.2*10^-3

[H+] = 3.125*10^-12 M

use:

pH = -log [H+]

= -log (3.125*10^-12)

= 11.5051

use:

pOH = -log [OH-]

= -log (3.2*10^-3)

= 2.4949

Answers:

[H+] = 3.1*10^-12

[OH-] = 3.2*10^-3

pH = 11.51

pOH = 2.49

B)

[OH-] = 2*[Sr(OH)2]

= 2* 4.9*10^-4 M

= 9.8*10^-4 M

use:

[H+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H+] = (1.0*10^-14)/[OH-]

[H+] = (1.0*10^-14)/9.8*10^-4

[H+] = 1.02*10^-11 M

use:

pH = -log [H+]

= -log (1.02*10^-11)

= 10.9912

use:

pOH = -log [OH-]

= -log (9.8*10^-4)

= 3.0088

Answers:

[H+] = 1.0*10^-11

[OH-] = 9.8*10^-4

pH = 10.99

pOH = 3.01

C)

[OH-] = [KOH]

= 8.7*10^-5 M

use:

[H+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H+] = (1.0*10^-14)/[OH-]

[H+] = (1.0*10^-14)/8.7*10^-5

[H+] = 1.149*10^-10 M

use:

pH = -log [H+]

= -log (1.149*10^-10)

= 9.9395

use:

pOH = -log [OH-]

= -log (8.7*10^-5)

= 4.0605

Answers:

[H+] = 1.1*10^-10

[OH-] = 8.7*10^-5

pH = 9.94

pOH = 4.06